OpenStudy (anonymous):

cos(sin^-1(5/13)-tan^-1(3/4) step by step please

OpenStudy (anonymous):

What exactly are we doing?

OpenStudy (anonymous):

Trying to find the value, or verifying it?

Miracrown (miracrown):

\[\cos (\sin ^{-1} (\frac{ 5 }{ 13 }) )- \tan ^{^{-1}} (\frac{ 3 }{ 4 })\]

OpenStudy (anonymous):

cos = 12/13....

Miracrown (miracrown):

these actually will give different types of responses, as in in the end we will have a length minus degrees, is it possible tan^-1 is in another function as well?

OpenStudy (anonymous):

12/13 - arctan(3/4)

Miracrown (miracrown):

we will do the cos(sin^-1(5/13)) to start

Miracrown (miracrown):



Miracrown (miracrown):

now do you remember which sides of a triangle sin of an angle equal?

OpenStudy (anonymous):

i need the exact value of it step by step though because i am very confused. please!!

Miracrown (miracrown):

I am explaining this to you step-by-step, you have to play your contribution as well, I can't just solve it for you - you won't learn this way. After all, this is your problem not mine.

OpenStudy (anonymous):

i know, and i am trying to understand but i am very confused.

OpenStudy (anonymous):

??